OpenStudy (anonymous):

Solve factoring 8x^2+20x=12

pooja195 (pooja195):

Hi Welcome to opentudy :) Any ideas on what we should do first?

OpenStudy (anonymous):

I already divide by common number and turned it in to binomials But I don't know what to do next

pooja195 (pooja195):

Ok show me what you have

OpenStudy (anonymous):

I have gotten up to (2x-1)(x+3)

pooja195 (pooja195):

Ok have you heard of the Zero Product property?

OpenStudy (anonymous):

Yes

pooja195 (pooja195):

We would apply that here can you show me what 2 equations we would have when we use that property :)

OpenStudy (anonymous):

Yes

pooja195 (pooja195):

Ok show me :-)

OpenStudy (anonymous):

(2x-1)(x+3)=0

pooja195 (pooja195):

Correct :) so \[\huge~\rm~\bf~2x−1=0 ~or~ x+3=0 \] Solve both equations for x

OpenStudy (anonymous):

X=1/2,x=-3

pooja195 (pooja195):

Good and those would be our answers

OpenStudy (anonymous):

Omg thank u so much

pooja195 (pooja195):

yw :)

pooja195 (pooja195):

\(~~~~~~~~~~~~~~~~~~~~~~~~~~\Huge{\color{red}{\heartsuit}\color{orange}{\bigstar}\color{yellow}{\heartsuit}\color{green}{\bigstar}\color{blue}{\heartsuit}\color{purple}{\bigstar}\color{hotpink}{\heartsuit}}\)\(\Huge\bf\color{#FF0000}W\color{#FF4900}e\color{#FF9200}l\color{#FFDB00}c\color{#FFff00}o\color{#B6ff00}m\color{#6Dff00}e\color{#24ff00}~\color{#00ff00}t\color{#00ff49}o\color{#00ff92}~\color{#00ffDB}O\color{#00ffff}p\color{#00DBff}e\color{#0092ff}n\color{#0049ff}S\color{#0000ff}t\color{#2400ff}u\color{#6D00ff}d\color{#B600ff}y\color{#FF00ff}!\)\(~~~~~~~~~~~~~~~~~~~~~~~~~~\Huge{\color{red}{\heartsuit}\color{orange}{\bigstar}\color{yellow}{\heartsuit}\color{green}{\bigstar}\color{blue}{\heartsuit}\color{purple}{\bigstar}\color{hotpink}{\heartsuit}}\)\(\Large\bf\color{Lime}{Please,\ don't\ hesitate\ to\ ask\ questions!}\)\(\Large\bf\color{magenta}{We're\ here\ to\ help!}\Large\color{magenta}{\ddot\smile}\)

OpenStudy (anonymous):

X=1/2,x=-3

OpenStudy (anonymous):

(2x-1)(x+3)=0

OpenStudy (anonymous):

Yes

pooja195 (pooja195):

o.O

OpenStudy (wolf1728):

So, 8x^2+20x -12 = (2x-1)(x+3) ?

pooja195 (pooja195):

@wolf1728 we already solve for x.

OpenStudy (usukidoll):

changing to the quadratic form would've been easier and then factor out what's common in the 8, 20, and 12

pooja195 (pooja195):

I assumed they wanted to solve by factoring because of their post title.

OpenStudy (anonymous):

Ya the question was to solve by factoring

OpenStudy (anonymous):

Can someone help me on my new question I just posted

OpenStudy (usukidoll):

that equation wasn't in standard quadratic form though. so subtracting 12 from both sides yields \[8x^2+20x-12=0\] and then 4 is what 8,20, and 12 have in common so factor 4 out \[4(2x^2+5x-3)=0\] \[4(2x-1)(x+3)=0\] and then solve for x on each situation which is 1/2 and -3.

pooja195 (pooja195):

@UsukiDoll we did all of that ._.

OpenStudy (usukidoll):

standard quadratic form is ax^2+bx+c=0 ok then.. just showing that there is more than one way of solving this problem as long as none of the math rules are broken.

pooja195 (pooja195):

It's nice to know but the question had asked to solve by factoring

OpenStudy (anonymous):

Ok now can someone Plzz help me with my new question I just posted

OpenStudy (anonymous):

Can someone help me on my new question I just posted